# Collinear Altermagnets and their Landau Theories

### v1

@@ I Introduction
In ideal altermagnets, these spin symmetries impose a compensated collinear antiferromagnetic order (magnetization $\mathbf{M}\!=\!0$) and preserve spin as a good quantum number while lifting the spin-degeneracy often associated with Néel antiferromagnets [10, 4, 11]. Overall compensation is preserved through the symmetry-imposed constraint that constant energy surfaces in momentum space (and thus occupied electronic bands) display alternating spin patterns [12, 5, 13, 14, 9, 10, 15, 4, 7, 16]. These spin splittings are even under inversion regardless of whether the crystal is centrosymmetric [13, 14, 15, 4, 7, 16] and can follow $d$-wave, $g$-wave or $i$-wave form factors. These anisotropic spin-splitting patterns are directly tied to their ability to produce spin currents [17] and are related to underlying secondary multipolar order parameters [18].
In the presence of weak SOC, some altermagnets produce a large anomalous Hall response that does not arise from canting of their magnetic moments (i.e. weak ferromagnetism) [19, 8, 20, 21, 22]. Other altermagnets exhibit a wide range of novel responses brought to light in Refs. [7, 8, 23] including the thermal Hall effect [24], piezomagnetism [25], and anisotropic magnetoresistance [26, 27] among others, leading to a great deal of interest in the unconventional transport properties arising from altermagnetism. A crucial recent development has been direct experimental imaging of the altermagnetic spin splitting in candidate altermagnets MnTe and CrSb using photoemission spectroscopy both with and without spin polarization  [28, 29]. As the definition of altermagnetism is grounded in symmetry, it has implications for all magnetic degrees of freedom meaning that a characteristic pattern of spin splitting of electronic bands should coincide with an analogous chirality splitting pattern in the spin-wave spectrum. Evidence of such a splitting of the magnon bands has been found in MnTe using inelastic neutron scattering [30], but has not been observed in the insulating candidate MnF2 [31].
@@ II Altermagnets from their symmetries
We begin this section with a general review of altermagnetism, translating the essential ideas into the language of representation theory. We then use this reformulation to perform a complete classification of crystal symmetries that are compatible with altermagnetism.
@@
Based on these transformation properties of $\mathbf{N}$, one can frame the search for altermagnetic orders as identifying structures in which $\mathbf{N}$ transforms as a nontrivial 1D inversion-even irrep of the crystal point group. Practically, this can be accomplished by constructing a collinear antiferromagnetic (AFM) order on each Wyckoff position (WP) in each space group and isolating the cases where the corresponding irrep $\Gamma_{\mathbf{N}}$ under which $\mathbf{N}$ transforms obeys the symmetry constraints described above. A detailed procedure for accomplishing this for an arbitrary space group and WP is provided in Appendix C, and the complete set of WPs compatible with altermagnetism is given in Table 12. As there is a growing need for materials identification and design [47], these results may help focus search efforts.
@@
We note that in non-centrosymmetric groups, all nontrivial, real, one-dimensional irreps correspond to altermagnetism. If the magnetic unit cell is not enlarged, any collinear antiferromagnet in these space groups will necessarily be altermagnetic.
@@ III.2 Secondary Multipolar Order Parameters for zero SOC
Secondary multipolar order parameters have significant implications for the spin-splitting structure of electronic bands, and they determine entire classes of observable quantities that couple to $\mathbf{N}$ in the presence of SOC [18]. Momentum space multipoles have been utilized to classify spin-splitting [16] in altermagnets and in the broader context of electronic band structures in magnetic materials [7, 49, 50, 51]. Our results differ from these in that they fully exhaust all possibilities for collinear $\bm{Q}=\bf{0}$ altermagnets, and are applicable beyond the analysis of electronic spin-splitting.
@@ IV.2 Coupling to the Néel Vector at Finite SOC
Linear coupling between $\mathbf{N}$ and $\xi$ is allowed if their representations share at least one common irrep in their decompositions. This criterion is equivalent to the trivial irrep appearing in the decomposition of $\Gamma_{\xi}\otimes(aeV\otimes\Gamma_{\mathbf{N}})$. Recall that $aeV$ is the time reversal odd axial vector representation. We may recast this condition into a more practical form: that $\Gamma_{\mathbf{N}}$ must appear in the decomposition of $\Gamma_{\xi}\otimes aeV$.
@@
Table 2: The representation $\Gamma_{\xi}$ for quantities $\xi$ is guaranteed to couple linearly to $\mathbf{N}$ in the presence of SOC, based on the rank-$n$ multipole in the SOC-free limit. The representations are denoted by Jahn symbols, where $aV$ is a time-reversal odd ($a$) polar vector ($V$), $aeV$ is a time-reversal odd ($a$) axial vector ($eV$), $aeV^{2}$ is a time-reversal odd axial tensor of rank 2, and $aeV[V^{2}]$ is a time-reversal odd axial rank-3 axial tensor that is symmetric in two indices. The $A=5$ case is absent because this multipole is not minimal for any irrep of any point group (see Table 13).
@@ V.4 Point group $\mathbf{\overline{3}m}$
In the point group $\overline{3}m,$ it has been suggested that the insulating collinear antiferromagnetic state of hematite Fe2O3 below the Morin temperature $T_{M}=265K$ [84] is altermagnetic [15]. Magnetism in hematite has been a longstanding and ongoing topic of research [85, 86, 87, 88, 89, 90]. Proposed altermagnetic features of hematite have been investigated in Ref. [91]. Here, we develop the SO-free and SOC Landau theories for hematite and compare them with known material properties.
@@ VI Discussion
Further, the nature of non-centrosymmetric altermagnets has received limited attention. Due to the emergent inversion-symmetry of the band structure, there is a discrepancy between the lowest order allowed SO-free multipolar order parameter and the spin-splitting pattern in reciprocal space. It may be worth exploring different properties that would inherit the lowest-order multipolar symmetry.
@@ Appendix A Circumventing Spin Groups
In the Landau theory, conventionally, one uses the symmetry group of the ordered phase, in which the (primary and secondary) order parameters transform as the fully symmetrical (trivial) irreducible representation. The symmetry groups of ideal altermagnetic phases correspond to spin groups [4, 15]. Spin point groups describing SO-free cases are subgroups of $\mathbf{O}^{s}(3)\times\mathbf{O}(3)$, where $\mathbf{O}^{s}(3)$ and $\mathbf{O}(3)$ contain the proper and improper rotations in spin space ($R_{s}$) and real space ($R_{l}$), respectively [37, 44]. The improper rotations in spin space contain the time-reversal operator (spin-inversion), $\tau$, and the improper rotations in real space include the inversion element, $I$. Group elements of $\mathbf{O}^{s}(3)\times\mathbf{O}(3)$ are usually written as $[R_{s}\|R_{l}]$, where the first element acts on spin and the second on latticed degrees of freedom [37, 44, 45].
@@ Appendix C Altermagnetic Structures Algorithm: Technical Details
The group elements of $\mathbf{G}$ can be expressed in Wigner-Seitz notation as $[f|\vec{t}]$, where $f\in\mathbf{F}$ is some $\mathbf{O}(3)$ matrix, and $\vec{t}$ is a three-dimensional translation vector666the single vertical bar distinguishes space group elements from the more general spin group notation [95, 94]. This group element transforms the atomic position $\vec{r}$ to $f\vec{r}+\vec{t}$.
@@ Appendix I Table of multipoles coupling to $\Gamma_{\mathbf{N}}$ in the SOC-free limit
$\mathbf{\frac{4}{m}mm}$ $A_{2g}$$\{xy(x^{2}-y^{2})\}$ $aeV[V^{2}]$ $zR_{z}\left(yN_{x}-xN_{y}\right),y^{2}N_{y}R_{x}-x^{2}N_{x}R_{y}$, $xy\left(N_{x}R_{x}-N_{y}R_{y}\right),x^{2}N_{y}R_{x}-y^{2}N_{x}R_{y}$, $zN_{z}\left(yR_{x}-xR_{y}\right),z^{2}\left(N_{y}R_{x}-N_{x}R_{y}\right)$

### v3

@@ I Introduction
In ideal altermagnets, these spin symmetries impose a compensated collinear antiferromagnetic order (magnetization $\mathbf{M}\!=\!0$) and preserve spin as a good quantum number while lifting the spin-degeneracy often associated with Néel antiferromagnets [10, 4, 11]. Overall compensation is preserved through the symmetry-imposed constraint that constant energy surfaces in momentum space (and thus occupied electronic bands) display alternating spin patterns [12, 5, 13, 14, 9, 10, 15, 4, 7, 16]. These spin splittings are even under inversion regardless of whether the crystal is centrosymmetric [13, 14, 15, 4, 7, 16] and can follow $d$-wave, $g$-wave or $i$-wave form factors. These anisotropic spin-splitting patterns are directly tied to their ability to produce spin currents [17, 18] and are related to underlying secondary multipolar order parameters [19].
In the presence of weak SOC, some altermagnets produce a large anomalous Hall response that does not arise from canting of their magnetic moments (i.e. weak ferromagnetism) [8, 20, 21, 22, 23]. Other altermagnets exhibit a wide range of novel responses brought to light in Refs. [7, 8, 24] including the thermal Hall effect [25], piezomagnetism [26], and anisotropic magnetoresistance [27, 28], and topological transitions [29] among others, leading to a great deal of interest in the unconventional transport properties arising from altermagnetism. A crucial recent development has been direct experimental imaging of the altermagnetic spin splitting in candidate altermagnets MnTe and CrSb using photoemission spectroscopy both with and without spin polarization  [30, 31]. As the definition of altermagnetism is grounded in symmetry, it has implications for all magnetic degrees of freedom meaning that a characteristic pattern of spin splitting of electronic bands should coincide with an analogous chirality splitting pattern in the spin-wave spectrum. Evidence of such a splitting of the magnon bands has been found in MnTe using inelastic neutron scattering [32], but has not been observed in the insulating candidate MnF2 [33].
@@ II Altermagnets from their symmetries
We begin this section with a general review of altermagnetism, translating the essential ideas into the language of representation theory. We then use this reformulation to perform a complete classification of crystal symmetries that are compatible with altermagnetism. We note that our analysis concerns altermagnetism arising from staggered dipolar order, and therefore does not encompass scenarios in which orbital ordering [35, 36], or ferromultipolar order with zero dipolar moment [37, 36], is responsible for altermagnetic spin-splitting.
@@
Based on these transformation properties of $\mathbf{N}$, one can frame the search for altermagnetic orders as identifying structures in which $\mathbf{N}$ transforms as a nontrivial 1D inversion-even irrep of the crystal point group. Practically, this can be accomplished by constructing a collinear antiferromagnetic (AFM) order on each Wyckoff position (WP) in each space group and isolating the cases where the corresponding irrep $\Gamma_{\mathbf{N}}$ under which $\mathbf{N}$ transforms obeys the symmetry constraints described above. A detailed procedure for accomplishing this for an arbitrary space group and WP is provided in Appendix C, and the complete set of WPs compatible with altermagnetism is given in Table 12. As there is a growing need for materials identification and design [52], these results may help focus search efforts. We note here that our analysis encompasses all viable WP (i.e. those of multiplicity two or greater), and is thus distinct from and expands upon the work of [53].
@@
We note that in non-centrosymmetric groups, all nontrivial, real, one-dimensional irreps correspond to symmetry-compensated collinear magnetic order, coinciding with altermagnetism. [^2] We do not address compensated ferrimagnetism, which is possible when relaxing the constraint that compensation is symmetry-enforced. If the magnetic unit cell is not enlarged, any collinear antiferromagnet in these space groups will necessarily be altermagnetic.
@@ III.2 Secondary Multipolar Order Parameters for zero SOC
Secondary multipolar order parameters have significant implications for the spin-splitting structure of electronic bands, and they determine entire classes of observable quantities that couple to $\mathbf{N}$ in the presence of SOC [19]. Momentum space multipoles have been utilized to classify spin-splitting [16, 53] and as order parameters [29] in altermagnets and in the broader context of electronic band structures in magnetic materials [7, 55, 56, 57]. Our results differ from these in that they fully exhaust all possibilities for collinear $\bm{Q}=\bf{0}$ altermagnets, and are applicable beyond the analysis of electronic spin-splitting.
@@ IV.2 Coupling to the Néel Vector at Finite SOC
Linear coupling between $\mathbf{N}$ and $\xi$ is allowed if their representations share at least one common irrep in their decompositions. This criterion is equivalent to the trivial irrep appearing in the decomposition of $\Gamma_{\xi}\otimes(aeV\otimes\Gamma_{\mathbf{N}})$. Recall that $aeV$ is the time reversal odd axial vector representation. We may recast this condition into a more practical form: that $\Gamma_{\mathbf{N}}$ must appear in the decomposition of $\Gamma_{\xi}\otimes aeV$. That this condition is equivalent may be seen by first invoking associativity of the direct product, so that we seek the trivial irrep in $(\Gamma_{\xi}\otimes aeV)\otimes\Gamma_{\mathbf{N}}.$ Then, it is clear that $(\Gamma_{\xi}\otimes aeV)$ must contain $\Gamma_{\mathbf{N}}$ in its decomposition for the trivial irrep to appear in this product.
@@
Table 2: The representation $\Gamma_{\xi}$ for quantities $\xi$ is guaranteed to couple linearly to $\mathbf{N}$ in the presence of SOC, based on the rank-$n$ multipole in the SOC-free limit. The representations are denoted by Jahn symbols, where $aV$ is a time-reversal odd ($a$) polar vector ($V$), $aeV$ is a time-reversal odd ($a$) axial vector ($eV$), $aeV^{2}$ is a time-reversal odd axial tensor of rank 2, and $aeV[V^{2}]$ is a time-reversal odd axial rank-3 axial tensor that is symmetric in two indices. The $n=5$ case is absent because this multipole is not minimal for any irrep of any point group (see Table 13).
@@ IV.5 Experimental Signatures of Altermagnetism
This result can be made considerably stronger: altermagnets whose lowest rank multipoles have $n>2$ have vanishing spin conductivity. To see this, note that the only axial, time-odd quantities that transform as a vector spin space that can be created using $\mathbf{N}$ are of the form $f(|\mathbf{N}|^{2})\mathbf{N}$ where $f$ is an arbitrary function. As $|\mathbf{N}|^{2}$ transforms trivially, this again transforms as $\Gamma_{\mathbf{N}}$ and so can only appear in $\bm{\sigma}^{s}_{\mu\nu}$ if $\Gamma_{\mathbf{N}}\subseteq[V^{2}]$. Thus altermagnets with $n=4$ or $n=6$ (i.e. with $g$-wave or $i$-wave spin splitting) do not have spin currents generated by electric fields in the SO-free limit.
We comment on the comparison between our framework and standard techniques at finite SOC using the black & white groups. One may just as well analyze all possible couplings at finite SOC; the result would be a potentially longer list of quantities than those listed in Table 5. However, this approach would not distinguish between properties arising from altermagnetism, and those simply arising from finite SOC. Herein lies the primary benefit of our framework: whereas the properties listed in Table 5 highlight only those quantities originating from ideal altermagnetism in the SO-free limit, our framework offers a bridge between the SO-free and SO-coupled theories. In addition, this perspective allows one to predict altermagnetic couplings solely on the basis of the lowest order pseudoprimary multipole.
There is a growing body of literature examining the connection between the Néel vector and various physical properties in altermagnets. For example, the anomalous Hall conductivity or magnetization in the presence of SOC has been explored in works including Ref. [69, 70, 27, 23, 8, 71, 19], among others. Here, we expand upon the previous literature by making a series of predictions applicable to all symmetry classes of tensors that may be seen as arising from altermagnetism, and by providing a coherent framework for understanding how these couplings arise.
@@ V.4 Point group $\mathbf{\overline{3}m}$
In the point group $\overline{3}m,$ it has been suggested that the insulating collinear antiferromagnetic state of hematite Fe2O3 below the Morin temperature $T_{M}=265K$ [92] is altermagnetic [15]. Magnetism in hematite has been a longstanding and ongoing topic of research [93, 94, 95, 96, 97, 98]. Proposed altermagnetic features of hematite have been investigated in Ref. [99], and recently, chiral splitting of magnons in hematite has been investigated [100]. Here, we develop the SO-free and SOC Landau theories for hematite and compare them with known material properties.
@@ VI Discussion
Further, the nature of non-centrosymmetric altermagnets has received limited attention [101]. Due to the emergent inversion-symmetry of the band structure, there is a discrepancy between the lowest order allowed SO-free multipolar order parameter and the spin-splitting pattern in reciprocal space. It may be worth exploring different properties that would inherit the lowest-order multipolar symmetry.
@@ Appendix A Circumventing Spin Groups
In the Landau theory, conventionally, one uses the symmetry group of the ordered phase, in which the (primary and secondary) order parameters transform as the fully symmetrical (trivial) irreducible representation. The symmetry groups of ideal altermagnetic phases correspond to spin groups [4, 15]. Spin point groups describing SO-free cases are subgroups of $\mathbf{O}^{s}(3)\times\mathbf{O}(3)$, where $\mathbf{O}^{s}(3)$ and $\mathbf{O}(3)$ contain the proper and improper rotations in spin space ($R_{s}$) and real space ($R_{l}$), respectively [42, 49]. The improper rotations in spin space contain the time-reversal operator (spin-inversion), $\tau$, and the improper rotations in real space include the inversion element, $I$. Group elements of $\mathbf{O}^{s}(3)\times\mathbf{O}(3)$ are usually written as $[R_{s}\|R_{l}]$, where the first element acts on spin and the second on lattice degrees of freedom [42, 49, 50].
@@ Appendix C Altermagnetic Structures Algorithm: Technical Details
The group elements of $\mathbf{G}$ can be expressed in Wigner-Seitz notation as $[f|\vec{t}]$, where $f\in\mathbf{F}$ is some $\mathbf{O}(3)$ matrix, and $\vec{t}$ is a three-dimensional translation vector777the single vertical bar distinguishes space group elements from the more general spin group notation [105, 104]. This group element transforms the atomic position $\vec{r}$ to $f\vec{r}+\vec{t}$.
@@ Appendix I Table of multipoles coupling to $\Gamma_{\mathbf{N}}$ in the SOC-free limit
$\mathbf{4/mmm}$ $A_{2g}$$\{xy(x^{2}-y^{2})\}$ $aeV[V^{2}]$ $zR_{z}\left(yN_{x}-xN_{y}\right),y^{2}N_{y}R_{x}-x^{2}N_{x}R_{y}$, $xy\left(N_{x}R_{x}-N_{y}R_{y}\right),x^{2}N_{y}R_{x}-y^{2}N_{x}R_{y}$, $zN_{z}\left(yR_{x}-xR_{y}\right),z^{2}\left(N_{y}R_{x}-N_{x}R_{y}\right)$